OpenStudy (anonymous):

A sample of sodium hydroxide (NaOH) has a mass of 160.0 g. The molar mass of NaOH is 40.00 g/mol. How many moles of NaOH does this sample contain?

ganeshie8 (ganeshie8):

simply divide ?

ganeshie8 (ganeshie8):

\[\large 160 g \times \dfrac{1}{40g/mol}\]

ganeshie8 (ganeshie8):

\[\large 160 g \times \dfrac{mol}{40g}\]

ganeshie8 (ganeshie8):

simplify

OpenStudy (lovelyharmonics):

hes an amazing helper. you see the blue button that says best response? you should click on that for him c:

OpenStudy (lovelyharmonics):

thats just common curtesy on open study

OpenStudy (anonymous):

hi my name is zahara could you help me with my question